Rene is 6 years older more than 2 times her younger sister Megan the sum of their ages is 33
Answer by stanbon(75887)

Rene is 6 years older more than 2 times her younger sister Megan the sum of their ages is 33
---------------------
R = 2M + 6
R + M = 33
-----------------------
Substitute for "R" and solve for "M"::
2M+6 + M = 33
--------------
3M = 27
----
Megan's age = 9 yrs
Rene"s age = 2M+6 = 24 yrs
-----------------
Cheers,
Stan H.
------------